# SARatrX: Towards Building A Foundation Model for SAR Target Recognition

### v2

@@ Abstract
Despite the remarkable progress in synthetic aperture radar automatic target recognition (SAR ATR), recent efforts have concentrated on the detection or classification of a specific and coarse category, e.g., vehicles, ships, airplanes, or buildings. One of the fundamental limitations of the top-performing SAR ATR methods is that the learning paradigm is supervised, task-specific, limited-category, closed-world learning, which depends on massive amounts of accurately annotated samples that are expensively labeled by expert SAR analysts and has limited generalization capability and scalability. In this work, we make the first attempt towards building a foundation model for SAR ATR, termed SARatrX. SARatrX learns generalizable representations via self-supervised learning (SSL) and provides a basis for label-efficient model adaptation to generic SAR target detection and classification tasks. Specifically, SARatrX is trained on 0.18 M unlabelled SAR target samples, which are curated by combining contemporary benchmarks and constitute the largest publicly available dataset till now. Considering the characteristics of SAR images, a backbone tailored for SAR ATR is carefully designed, and a two-step SSL method endowed with multi-scale gradient features was applied to ensure the feature diversity and model scalability of SARatrX. The capabilities of SARatrX are evaluated on classification under few-shot and robustness settings and detection across various categories and scenes, and impressive performance is achieved, often competitive with or even superior to prior fully supervised, semi-supervised, or self-supervised algorithms. Our SARatrX and the curated dataset are released at https://github.com/waterdisappear/SARatrX to foster research into foundation models for SAR ATR and SAR image interpretation.
@@
Synthetic Aperture Radar (SAR), Object Recognition, Object Detection, Foundation Model, Self-supervised learning, Deep Learning, Masked Image Modeling
https://github.com/waterdisappear/SARatrX
@@ I Introduction
Synthetic aperture radar (SAR) [1, 2, 3, 4], developed from electromagnetic scattering in microwave frequency bands, plays a crucial role in active Earth observations, functioning effectively across diverse weather conditions and lighting environments. With the rapid development of SAR imaging techniques, high-resolution SAR images can be accessed more easily than before, enabling even more research opportunities for SAR image intelligent interpretation. SAR automatic target recognition (ATR), aiming at automatically localizing and classifying objects of interest (e.g., vehicles, ships, airplanes, or buildings) in SAR images, is a longstanding, important yet challenging problem in SAR image intelligent interpretation [5, 6, 7, 8]. SAR ATR plays an essential role in civil and national defense applications such as modern airport management, disaster management, urban planning and infrastructure monitoring, military reconnaissance, and maritime surveillance. Therefore, it has become an active research area for several decades [9, 10, 11, 12, 13, 14]. In the past decade, deep learning has brought tremendous success for SAR ATR [15, 5, 16, 17, 8]. Despite the significant progress, there are at least the following fundamental challenges that will need to be addressed to further advance the field of SAR ATR.
Figure 1: The motivation and performance of SARatrX. (a) SAR ATR includes various imaging conditions, targets, scenes, and sensors. However, due to high costs, the datasets are often collected in specific settings for certain tasks. For example, MSTAR [18] is a ten-type vehicle target classification dataset in the X-band and grass scenarios, while SAR-Aircraft is a seven-type aircraft target detection dataset collected from three airports and a C-band satellite. Several specialized algorithms have been proposed for application to these datasets. However, the differing target characteristics, scene background information, and sensor parameters have complicated the generalization of existing algorithms. As such, this paper aims to develop a SAR ATR foundation model, a generalized method for conducting various tasks. (b) SARatrX performed well across 5 datasets with 8 settings. It was superior to existing SSL methods (BIDFC [19]) for target classification in the fine-grained vehicle MSTAR dataset [18] with a few-shot setting. In addition, it performed well under extended operating conditions (EOCs) [20] (i.e., imaging conditions with variable depression angle (EOCs-Depression), target configuration (EOCs-Config), and version (EOCs-Version)). SARatrX also demonstrated competitive object detection performance with existing supervised methods applied to various categories (SARDet-100K [21] and OGSOD [22]), as well as specific categories for ships (SSDD [23]) and aircraft (SAR-AIRcraft [24]). Our study shows the potential of a foundation model for generalized SAR ATR.
@@
Firstly, task-specific property. One of the fundamental limitations of the current SAR ATR methods [25, 26, 27, 28, 29, 30] is that one model is trained and evaluated on one specific task. The detection and classification of a specific coarse category (e.g., vehicles, ships, airplanes, or buildings in Fig. 1.) all require their own deep models. As a result, the task-specific properties of these deep models pose significant challenges for training new tasks or developing a comprehensive SAR ATR system since each task must be learned independently from the ground up, requiring vast amounts of labeled data. This results in computational inefficiency, lower accuracy, and inconsistent results between the different models. Secondly, heavy reliance on supervised learning. Recent progress in SAR ATR, while substantial, has been limited to supervised learning, which heavily depends on massive amounts of accurately annotated target samples that are expensively labeled by expert SAR analysts and have limited generalization capability and scalability. However, the scarcity of expert SAR analysts cannot meet such an exhaustive requirement, leaving vast amounts of SAR images unlabelled and unexploited. Thirdly, the ignorance of SAR image characteristics in model designs. The imaging characteristics of SAR imagery differ significantly from those of optical imagery, leading to a significant domain gap between natural and SAR images. This raises significant challenges when one intends to transfer prior knowledge from the natural image domain. Different strong prior knowledge of SAR imagery, including speckle noise, discrete target appearances, and the lack of geometry, texture, and contour cues, needs special consideration when designing backbone architectures. Most of the current mainstream backbones and methods designed on natural images are not suitable for the aforementioned information. Finally, underdeveloped open-source ecosystem. Due to data sensitivity, the open-source ecosystem across the entire field is underdeveloped, making it challenging to share code and data publicly. Currently, there are no large and representative benchmark datasets for SAR ATR. As a result, this locks the potential of recent deep learning techniques for SAR ATR and significantly slows down the development of this field.
@@
Recently, the remarkable success of foundation models (FMs) [31, 32, 33, 34] has led to a learning paradigm shift in artificial intelligence. Foundation models [35], pretrained on extensive data in a task-agnostic manner (generally via Self-supervised learning), can be flexibly adapted to a wide range of downstream tasks. Self-supervised learning (SSL) [36, 37, 7, 38] can be used to mitigate label inefficiencies by exploring supervision in the data directly, thereby reducing the reliance on expensive expert labeling while efficiently scaling the data and models. FMs shine in a broad range of areas, including natural language processing, computer vision, speech recognition, and medical image analysis. FMs have also been explored in remote sensing image understanding, but it is mostly limited to the evaluation of optical data. To our knowledge, the huge potential of FMs for SAR image interpretation remains completely locked.
@@
In our preliminary work, a novel SSL method for SAR imagery named SAR Joint-Embedding Predictive Architecture (SAR-JEPA) [39] was proposed and demonstrated promising results. Given the aforementioned discussion, to fully unlock the potential of LMs for SAR image interpretation, we present the first attempt toward building a foundation model for SAR ATR and propose SARatrX that learns generalizable representations via SSL and provides a basis for label-efficient model adaptation to generic SAR target detection and classification tasks. While conceptually simple, existing methods for building FMs in other domains cannot be just directly applied to SAR ATR. We are still facing significant challenges, which we will discuss below and address with SARatrX.
@@
Pre-training datasets must include diverse target categories and imaging conditions to accommodate various downstream tasks. However, SAR ATR lacks a large-scale dataset, such as ImageNet [40], and the most common MSTAR dataset only includes fine-grained vehicle categories that are not suitable for larger-scale pre-training. As such, several SSL methods combine different datasets. Specifically, SARDet-100K [21] incorporates 9 SAR target detection sets. With the number of open-source datasets used for SAR ATR increasing, most of the open-source datasets were integrated into this study as part of the pre-training. A total of 14 classification and detection datasets with different target categories and imaging conditions were included in this study as a new pre-training dataset, SARDet-180K, to explore the potential of FMs, as seen in Table I.
@@
Model backbones aim to achieve better spatial representations in remote sensing images, especially for small target signatures in large imagery. Transformers and convolutional neural networks are among the most common architectures used for these tasks. Transformers offer better spatial resolution without downsampling, while convolutional neural networks (CNNs) provide higher efficiency through convolution kernels. The performance of different architectures was tested in MSFA [21] with SAR detection tasks, the results of which have shown the Swin transformer backbone to be the most effective for remote sensing images. As such, HiViT [41] was selected for use in this study, as it has the advantages of a Swin transformer and can drop patches in Masked Image Modeling (MIM).
@@
Self-supervised learning is complicated by SAR image quality, which is negatively affected by speckle noise in coherent imaging. The coherent imaging resulting visual features are also not as distinct or rich as in natural RGB images. Contrastive learning [19, 42] uses data augmentation and preprocessing to reduce noise, while MIM [43, 39, 21] applies various target feature for guided signals to suppress noise. As such, the primary task of SAR SSL is the enhance the quality of feature learning and guide signals. For example, PGIL [6] leveraged a sub-frequency feature of complex SAR images to learn physics information, while our SAR-JEPA [39] applied multi-scale gradient ratios to solve for multiplicative speckle noise and capture target shapes. Furthermore, multi-stage training [21] from ImageNet to SAR diminished the effects of noise on model diversity, as seen in Fig. 4. Thus, we applied two-step pretraining from ImageNet to SAR to increase model diversity during pretraining with SAR images. Besides, multiscale gradient features were used as high-quality guide signals for MIM with SAR images.
@@
Evaluation tasks need to comprehensively evaluate the performance of a foundation model for different tasks and settings. The availability of 3 open-source target datasets was utilized by first constructing a fine-grained classification dataset, SAR-VSA, with 25 categories to evaluate the effectiveness of the proposed improvements. A comprehensive comparison was then performed between the proposed SARatrX and existing methods for public classification and detection tasks.
@@
SARatrX achieved superior performance in 5 datasets across 8 task settings, as shown in Fig. 1, which is competitive with prior methods on various SAR ATR tasks (few-shot vehicle classification, robust classification, detection with specific categories or various categories). We hope that this work could advance the development of the intersection of general SAR target recognition and FMs. The primary contributions of this study can be summarized as follows:
@@ 1. $\bullet$
    We present the first foundation model called SARatrX, which learns generalizable representations via SSL from large-scale unlabelled data and provides a basis for generic SAR target detection and classification tasks.
@@ 2. $\bullet$
    We systematically investigate a foundation model framework for SAR ATR. We build a currently the largest publicly available pre-training dataset SARdet-180k and fully discussed the model architecture and proposed SSL method with many comparisons.
@@ 3. $\bullet$
    The ability of SARatrX is comprehensively evaluated with various SAR ATR tasks, such as few-shot classification, robust classification, ship detection, aircraft detection, and detection with various categories tasks.
The remainder of this paper is organized as follows. Sec. II introduces related work in remote sensing and SAR ATR. Sec. III discusses the proposed foundation model (SARatrX). Secs. IV and V conduct extensive experiments to demonstrate the superiority of the proposed method. Sec. VI concludes the paper and discusses future work.
@@ II Related Work
Visual foundation models are actively being used for remote sensing applications and several different algorithms have recently been proposed for various modalities and tasks. However, this study focuses on a foundation model used for SAR ATR (i.e., SAR image-based target classification and object detection). In the following sections, we introduce recent developments in remote sensing foundation models.

### v3

# SARATR-X: Towards Building A Foundation Model for SAR Target Recognition

@@ Abstract
Despite the remarkable progress in synthetic aperture radar automatic target recognition (SAR ATR), recent efforts have concentrated on detecting and classifying a specific category, e.g., vehicles, ships, airplanes, or buildings. One of the fundamental limitations of the top-performing SAR ATR methods is that the learning paradigm is supervised, task-specific, limited-category, closed-world learning, which depends on massive amounts of accurately annotated samples that are expensively labeled by expert SAR analysts and have limited generalization capability and scalability. In this work, we make the first attempt towards building a foundation model for SAR ATR, termed SARATR-X. SARATR-X learns generalizable representations via self-supervised learning (SSL) and provides a cornerstone for label-efficient model adaptation to generic SAR target detection and classification tasks. Specifically, SARATR-X is trained on 0.18 M unlabelled SAR target samples, which are curated by combining contemporary benchmarks and constitute the largest publicly available dataset till now. Considering the characteristics of SAR images, a backbone tailored for SAR ATR is carefully designed, and a two-step SSL method endowed with multi-scale gradient features was applied to ensure the feature diversity and model scalability of SARATR-X. The capabilities of SARATR-X are evaluated on classification under few-shot and robustness settings and detection across various categories and scenes, and impressive performance is achieved, often competitive with or even superior to prior fully supervised, semi-supervised, or self-supervised algorithms. Our SARATR-X and the curated dataset are released at https://github.com/waterdisappear/SARATR-X to foster research into foundation models for SAR image interpretation.
@@
Synthetic Aperture Radar, Target Recognition, Object Detection, Foundation Model, Self-supervised learning, Deep Learning, Masked Image Modeling
@@ I Introduction
Synthetic aperture radar (SAR) [1, 2, 3, 4], developed from electromagnetic scattering in microwave bands, plays a crucial role in active Earth observations, functioning effectively across diverse weather conditions and lighting environments. With the rapid development of SAR imaging techniques, high-resolution SAR images can be accessed more easily than before, enabling even more research opportunities for intelligent interpretation of SAR images. SAR automatic target recognition (ATR), aiming at automatically localizing and classifying objects of interest (e.g., vehicles, ships, airplanes, or buildings) in SAR images, is a longstanding, important yet challenging problem in SAR image intelligent interpretation [5, 6, 7, 8]. SAR ATR [^1] In this paper, we focus on SAR ATR’s classification and detection tasks. The classification aims to recognize the class and type of a single object, and the detection identifies the class and location of a single or multiple objects. plays an essential role in civil and national defense applications such as modern airport management, disaster management, urban planning and infrastructure monitoring, military reconnaissance, and maritime surveillance. Therefore, it has become an active research area for several decades [9, 10, 11, 12, 13, 14]. In the past decade, deep learning has brought tremendous success for SAR ATR [15, 16, 17]. Despite the significant progress, the following fundamental challenges will need to be addressed to advance the field of SAR ATR.
@@
Firstly, task-specific property. One of the fundamental limitations of the current ATR methods [18, 19, 20, 21, 22] is that one model is trained and evaluated on one specific task. The detection and classification of a specific coarse category (e.g., vehicles, ships, airplanes, or buildings in Fig. 1.) all require their own deep models. As a result, the task-specific properties of these deep models pose significant challenges for training new tasks or developing a comprehensive SAR ATR system since each task must be learned independently from the ground up, requiring vast amounts of labeled data. This results in computational inefficiency, lower accuracy, and inconsistent results between the different models. Secondly, heavy reliance on supervised learning. Recent progress in SAR ATR [23, 24, 25], while substantial, has been limited to supervised learning, which heavily depends on massive amounts of accurately annotated target samples that are expensively labeled by expert SAR analysts and have limited generalization capability and scalability. However, the scarcity of expert SAR analysts cannot meet such an exhaustive requirement, leaving vast amounts of SAR images unlabelled and unexploited. Thirdly, the ignorance of SAR image characteristics in model designs. The imaging characteristics of SAR imagery differ significantly from those of optical imagery, leading to a significant domain gap between natural and SAR images. This raises significant challenges when one intends to transfer prior knowledge from the natural image domain. Different strong prior knowledge of SAR imagery, including speckle noise, discrete target appearances, and the lack of geometry, texture, and contour cues, needs special consideration when designing backbone architectures and learning strategies. Most of the current mainstream backbones and methods designed on natural images are not suitable for the aforementioned information. Finally, underdeveloped open-source ecosystem. Due to data sensitivity, the open-source ecosystem across the entire field is underdeveloped, making it challenging to share code and data publicly. Currently, there are no large and representative benchmark datasets for SAR ATR. As a result, this locks the potential of recent deep learning techniques for SAR ATR and significantly slows down the development of this field.
@@
Recently, the remarkable success of foundation models (FMs) [26, 27, 28, 29] has led to a learning paradigm shift in artificial intelligence. Foundation models [30], pretrained on extensive data in a task-agnostic manner (generally via self-supervised learning), can be flexibly adapted to a wide range of downstream tasks. Self-supervised learning (SSL) [31, 32, 33, 34, 35] can be used to mitigate label inefficiencies by exploring supervision in the data directly, thereby reducing the reliance on expensive expert labeling while efficiently scaling the data and models. FMs shine in a broad range of areas, including natural language processing, computer vision, speech recognition, and medical image analysis. As summarized in Table I, FMs have also been explored in remote sensing image understanding, but they are mostly limited to the evaluation of optical data. To our knowledge, the huge potential of FMs for SAR image interpretation remains completely locked.
Figure 1: Various specialized SAR ATR datasets and tasks. SAR ATR includes various imaging conditions (i.e. operating condition), such as targets, scenes, and sensors. However, the datasets are often collected in specific settings for certain tasks due to high costs. For example, MSTAR [36] is a ten-type vehicle target classification dataset in the X-band and grass scenarios, and SAR-Aircraft is a seven-type aircraft detection dataset collected from three airports and a C-band satellite. Specialized algorithms have been proposed for these datasets. However, the differing target characteristics, scene information, and sensor parameters have complicated the generalization of existing algorithms. As such, this paper aims to develop a SAR ATR foundation model, a generalized method for conducting various tasks.
@@
In our preliminary work, a novel SSL method for SAR imagery named SAR Joint-Embedding Predictive Architecture (SAR-JEPA) [37] was proposed and demonstrated promising results. Given the aforementioned discussion, to fully unlock the potential of FMs for SAR image interpretation, we present the first attempt toward building a foundation model for SAR ATR and propose SARATR-X that learns generalizable representations via SSL and provides a basis for label-efficient model adaptation to generic SAR target detection and classification tasks. While conceptually simple, existing methods for building FMs in other domains cannot be just directly applied to SAR ATR. We are still facing significant challenges, which are discussed and addressed below.
@@
Pre-training datasets must include diverse target categories and imaging conditions to accommodate various downstream tasks. However, SAR ATR lacks a large-scale dataset, such as ImageNet [38], and the most used MSTAR dataset only includes fine-grained vehicle categories that are not suitable for larger-scale pre-training (see Table II). As such, SARDet-100K [39] incorporates 9 SAR target detection datasets. With the number of open-source SAR ATR datasets increasing, most of the datasets were integrated into this study as part of the pre-training. A total of 14 classification and detection datasets with different target categories and imaging conditions were included as a new pre-training dataset, SARDet-180K, to explore the potential of FMs, as seen in Table III.
@@
Model backbones aim to achieve better spatial resolution representations in remote sensing images, especially for small targets in large imagery. Transformers and convolutional neural networks are among the most common architectures for these tasks. As shown in Table I, the Transformer offers better spatial resolution without downsampling, which is most commonly used in recent studies. As such, HiViT [40] was selected as it has the advantages of a Swin transformer high-resolution input and can drop patches in masked image modeling (MIM).
@@
Self-supervised learning is complicated by SAR image quality, which is negatively affected by speckle noise in coherent imaging. The coherent imaging resulting visual features are also not as distinct or rich as in nature RGB images. Contrastive learning [41, 42] uses data augmentation and preprocessing to reduce noise, while MIM [43, 37, 39] applies various target feature for guided signals to suppress noise (see Table II). As such, the primary task of SAR SSL is to enhance the quality of feature learning and guide signals. For example, PGIL [6] leveraged a sub-frequency feature of complex SAR images to learn physics information, while our SAR-JEPA [37] applied multi-scale gradient ratios to solve for multiplicative speckle noise and capture target shapes. Furthermore, multi-stage training [39] from ImageNet to SAR diminished the effects of noise on model diversity, as seen in Fig. 5. Thus, we applied two-step pre-training from ImageNet to SAR to increase model diversity during pre-training with SAR images. Besides, multi-scale gradient features were used as high-quality guide signals for MIM with SAR images.
@@
Evaluation tasks need to comprehensively evaluate the performance of a foundation model for different tasks and settings. Three open-source target datasets were utilized by first constructing a fine-grained classification dataset, SAR-VSA, with 25 categories to evaluate the effectiveness of the proposed improvements. A comprehensive comparison was then performed between the proposed SARATR-X and existing methods for public classification and detection tasks.
@@
SARATR-X achieved superior performance in 5 datasets across 8 task settings, as shown in Fig. 2, which is competitive with prior methods on various SAR ATR tasks (classification with few-shot and robustness setting and detection with specific categories or various categories). We hope that this work could advance the development of the intersection of general SAR target recognition and foundation models.
The primary contributions of this study can be summarized as follows:
Figure 2: Results on classification and detection tasks. SARATR-X performed well across 5 datasets with 8 settings. It was superior to existing SSL methods (BIDFC [41]) for target classification in the fine-grained vehicle MSTAR dataset [36] with a few-shot setting. In addition, it performed well under extended operating conditions (EOCs) [44] (i.e., imaging conditions with variable depression angle (EOCs-Depression), target configuration (EOCs-Config), and version (EOCs-Version)). SARATR-X also demonstrated competitive object detection performance with existing supervised methods applied to various categories (SARDet-100K [39] and OGSOD [45]), as well as specific categories for ships (SSDD [46]) and aircraft (SAR-AIRcraft [47]). Our study shows the potential of a foundation model for SAR ATR.
@@ 1. $\bullet$
    We present the first foundation model called SARATR-X, which learns generalizable representations via SSL from large-scale unlabelled data and provides a cornerstone for generic SAR target detection and classification tasks.
@@ 2. $\bullet$
    We systematically investigate a foundation model framework for SAR ATR. We build the largest publicly available pre-training dataset, SARdet-180k, and fully discuss the model architecture and proposed SSL method with many comparisons.
@@ 3. $\bullet$
    SARATR-X is evaluated comprehensively with various SAR ATR tasks, such as few-shot classification, robust classification, ship detection, aircraft detection, and detection with various categories.
The remainder of this paper is organized as follows. Sec. II introduces related work in remote sensing and SAR ATR. Sec. III discusses the proposed foundation model (SARATR-X). Secs. IV and V conduct extensive experiments to demonstrate the superiority of the proposed method. Sec. VI concludes the paper and discusses future work.
@@ II Related Work
TABLE I: A list of visual FMs used for remote sensing. There are many FMs are available for various modalities and tasks, but existing research (FG-MAE, SkySense, and OFA-Net) has focused on scene-level tasks with the SAR modality. SARATR-X represents an extension of our previous work (SAR-JEPA and MSFA [37, 39]) and advances the state of FMs for SAR ATR. SSL methods can be divided into two categories: contrastive (i.e., contrastive learning used to obtain invariant features) and generative (i.e., masked image modeling used to generate and predict features). SD: stable diffusion. Swin: Swin transformer.
Foundation Model Year Modality Dataset Backbone SSL Tasks
SatMAE [48] 2022 Multi-spectral, RGB fMoW RGB/Sentinel ViT Generative Scene classification and semantic segmentation
RVSA [49] 2022 RGB MillionAID ViT, ViTAE Generative Scene classification, object detection, and semantic segmentation
RingMo [50] 2022 RGB Self-built ViT, Swin Generative Scene classification, object detection, semantic segmentation, and change detectionn
RingMo-Sense [51] 2023 RGB Self-built Video Swin Generative object detection and tracking, images segmentation, and 3 prediction tasks
CMID [52] 2023 RGB MillionAID ResNet, Swin Generative & Contrastive Scene classification, object detection, and semantic segmentation
GFM [53] 2023 RGB GeoPile Swin Generative Scene classification, semantic segmentation, change detection, and super-resolution
DiffusionSat [54] 2023 RGB fMoW, Satlas, SpaceNet SD Generative Image generation, super-resolution, temporal generation, and in-painting
Scale-MAE [55] 2023 RGB fMoW RBG ViT Generative Scene classification, object detection, and semantic segmentation
FG-MAE [43] 2023 Multi-spectral, SAR SSL4EO-S12 ViT Generative Scene classification and semantic segmentation
SMLFR [56] 2024 RGB GeoSense ConvNeXt Generative Object detection, semantic segmentation, and change detection
SkySense [57] 2024 3 modalities HSROIs, TMsI, TSARI Swin, ViT Contrastive Scene classification, object detection, semantic segmentation, and change detection
OFA-Net [58] 2024 4 modalities 5 multi-modalities datasets ViT Generative Scene Classification and semantic segmentation
SARATR-X 2024 SAR ImageNet & 14 SAR datasets HiViT Generative Target classification and object detection
TABLE II: Related SSL work in SAR ATR. They has involved the use of various datasets, architectures, and SSL algorithms. Inspired by previous work, this study presents the first SAR ATR foundation model, SARATR-X.
Method Year Dataset Backbone SSL Tasks Description
RotANet [59] 2021 MSTAR CNN Generative Target classsification (MSTAR) Predicting rotational patterns of targets as a loss regularization term for classification task.
UACL [60] 2021 MSTAR, FUSAR-Ship CNN Contrastive Target classsification (MSTAR) Contrastive defense method to enhance model robustness to different adversarial attack.
PGIL [6] 2022 Sea-ice, Urban ResNet-18 Contrastive Scene classsification (sea-ice, urban) Contrastive learning between sub-frequency features of complex images and deep features of amplitude images.
BIDFC [41] 2022 MSTAR, OpenSARShip ResNet-18 Contrastive Target classsification (MSTAR, OpenSARShip) Weakly contrastive learning for pre-training in fine-grained datasets with similar sample.
TSCL [42] 2023 MSTAR ResNet-50 Contrastive Target classsification (MSTAR, OpenSARShip) Pre-processing of SAR images for noise removal to improve image quality for Contrastive learning.
FG-MAE [43] 2023 SSL4EO-S12 ViT-S Generative Scene classification (EuroSAT, BigEarthNet-MM), semantic segmentation (DFC2020) Hand-craft feature HOG as the target signal for MIM in SAR images.
SAR-JEPA [37] 2024 MSAR, SAR-Ship, SARSim, SAMPLE ViT-B Generative Target classification (MSTAR, FUSARShip, SAR-ACD) Local reconstruction and multi-scale gradient feature for MIM with small object in noise SAR images.
MSFA [39] 2024 ImageNet, DOTA, SARDet-100k ConvNeXt, Swin-B, Van Generative Object detection (SARDet-100K, SSDD, HRSID) Multi-stage pre-training strategy for MIM from RGB images to SAR images.
SARATR-X 2024 ImageNet, 14 SAR target datasets HiViT-B Generative Target classification and object detection (5 datasets and 8 settings) A first foundation model for target recognition in SAR Images.
Visual foundation models are actively being used for remote sensing applications, and different algorithms have recently been proposed for various modalities and tasks. This study focuses on a foundation model used for SAR ATR (i.e., SAR image-based target classification and object detection). In the following sections, we introduce recent developments in remote sensing and SAR foundation models.